# LLMs Love Python: A Study of LLMs’ Bias for Programming Languages and Libraries

### v1

@@ Abstract.
Programming language and library choices are crucial to software reliability and security. Poor or inconsistent choices can lead to increased technical debt, security vulnerabilities, and even catastrophic failures in safety-critical systems. As Large Language Models (LLMs) play an increasing role in code generation, it is essential to understand how they make these decisions. However, little is known about their preferences when selecting programming languages and libraries for different coding tasks. To fill this gap, this study provides the first in-depth investigation into LLM preferences for programming languages and libraries used when generating code. We assess the preferences of eight diverse LLMs by prompting them to complete various coding tasks, including widely-studied benchmarks and the more practical task of generating the initial structural code for new projects (a crucial step that often determines a project’s language or library choices).
Our findings reveal that LLMs heavily favour Python when solving language-agnostic problems, using it in 90%-97% of cases for benchmark tasks. Even when generating initial project code where Python is not a suitable language, it remains the most-used language in 58% of instances. Moreover, LLMs contradict their own language recommendations in 83% of project initialisation tasks, raising concerns about their reliability in guiding language selection. Similar biases toward well-established libraries further create serious discoverability challenges for newer open-source projects. These results highlight the need to improve LLMs’ adaptability to diverse programming contexts and to develop mechanisms for mitigating programming language and library bias.
@@ 1. Introduction
The last few years have seen Large Language Models (LLMs) make huge advances (naveedComprehensiveOverviewLarge2024a), particularly excelling in code generation tasks (jiangSurveyLargeLanguage2024). Widely accessible tools like ChatGPT (openaiGPT4TechnicalReport2024) and GitHub Copilot (chenEvaluatingLargeLanguage2021; GitHubCopilotYour2024) enable developers of all experience levels to leverage generative AI in their workflow, leading to a significant impact on newly produced code (staffSurveyAIWave2024; lanzStabilityAICEO2023; mckenna25GooglesCode2024).
While LLMs have demonstrated impressive performance, little is known about their preferences when selecting programming languages and libraries for various coding tasks. Existing research has primarily focused on evaluating and improving their correctness, security, and efficiency in generating code (chenSurveyEvaluatingLarge2024), but the decision-making process behind language and library selection remains unexplored. This gap is particularly significant as LLMs are increasingly used in real-world software development workflows, where informed programming language and library choices are vital for ensuring project success. If these LLMs exhibit systematic biases or inconsistent recommendations, they may mislead developers into making flawed decisions, potentially compromising software reliability, security, and maintainability. In critical applications—such as healthcare, finance, and autonomous systems—these biases could lead to severe safety risks, security vulnerabilities, and system failures, endangering both users and infrastructure.
LLMs may develop strong preferences for certain programming languages or libraries due to various factors, such as the distribution of languages in their training data (wangExploringMultiLingualBias2024), the prevalence of certain libraries in open-source repositories, or biases introduced during model training and fine-tuning (zieglerFineTuningLanguageModels2020). For example, if an LLM is trained on codebases where a particular language dominates, it may exhibit a strong inclination toward generating solutions in that language, even when alternative languages would be more appropriate for the task. As LLMs are increasingly integrated into software development workflows, understanding these preferences is essential not only for assessing their reliability and applicability across diverse programming scenarios but also for developing strategies to mitigate potential biases and improve their adaptability.
To fill this gap, this paper provides the first empirical investigation into the programming language and library preferences of LLMs when writing code. As the first paper to consider and discuss such preferences, we focus on two primary aspects of coding: programming languages, as they are intrinsic to coding; and libraries, as they provide essential pre-built functionality, and play a crucial role in modern software development (somervilleSoftwareEngineeringGlobal2016).
@@
We measure LLM preferences by using prompts that require code generation without explicitly specifying the programming language or library — an approach that reflects common real-world scenarios, particularly for low-code users and developers initiating new projects (rasnayakaEmpiricalStudyUsage2024). The responses are then processed, with languages and libraries extracted for further analysis. To ensure a broad understanding of LLM preferences, eight widely-used LLMs are chosen for the study, covering both open- and closed-source and a range of sizes: GPT-4o (openaiGPT4TechnicalReport2024), GPT-3.5 (andrewpengGPT35TurboFinetuning), Claude3.5 (Haiku and Sonnet) (anthropicClaude3Model2024), Llama3.2 (grattafioriLlama3Herd2024), Qwen2.5 (huiQwen25CoderTechnicalReport2024), DeepSeekLLM (deepseek-aiDeepSeekLLMScaling2024) and Mistral7b (jiangMistral7B2023).
@@
For both programming languages and libraries, we examine LLM preferences in two key scenarios: (1) solving problems from widely-studied benchmarks, and (2) generating the initial code for a new project, a crucial phase where foundational technology choices—such as programming language and library selection—are made. Additionally, we examine whether all LLMs share similar preferences. We also assess their internal consistency by checking if the languages and libraries they recommend match those they actually use when generating code for the same task.
@@
Our experiments lead to the following primary observations: 1) All LLMs we study exhibit a significant bias toward Python as their preferred programming language. To our surprise, for scenarios where Python is not a suitable language, Python remains the dominant choice as the most-used language in 58% of cases. 2) For library usage, LLMs heavily favour top-ranked libraries such as NumPy, yet in up to 48% of cases, this usage is unnecessary and differs from the ground-truth solutions. 3) There is extremely low consistency between the programming languages and libraries LLMs recommend and those that they actually adopt, and the strong biases towards languages and libraries are universal across all the LLMs we study.
These findings reveal that widely used LLMs have notable diversity problems in their coding preferences, frequently straying from best practices and domain requirements. Moreover, such built-in biases may lead to homogenised code that limits creativity and niche solutions, and hinders the discoverability and growth of open-source software, which may ultimately weaken the effectiveness of LLM-based code generation.
The primary contributions of this paper are:
@@
  • Presenting the first empirical study into LLM preferences when generating code, revealing a clear lack of diversity across LLMs when choosing programming languages and libraries, with significant preferences towards Python and established libraries.
@@
  • Raising awareness of the need to measure and understand the diversity and preferences within LLM-generated code, providing a discussion of how this could impact the coding landscape and open-source software.
@@
  • Releasing our code, datasets and full results publicly, to encourage further investigation in this area:
    https://github.com/itsluketwist/llms-love-python
@@
The rest of the paper is organised as follows: Section 2 summarises work in related areas. Section 3 details the research questions, and the experiments conducted to answer them. Section 4 presents the results and answers to the research questions. Section 5 further discusses the findings and their broader impacts. Section 6 concludes the paper.
@@ 2. Related Work
To the best of our knowledge, this is the first paper to directly examine LLM preferences in programming languages and libraries for code generation. While no prior work specifically addresses this topic, several related problems have been explored in the literature.
@@
Bias and Unfairness in LLMs: Whether an LLM has preferences for specific programming languages and libraries is related to bias and whether an LLM provides an unfair representation of the world. Bias within an LLMs natural language responses has been extensively studied (gallegosBiasFairnessLarge2024), but bias within any generated code is widely under explored. Most works focus on trustworthiness of the code (patelEvaluatingInContextLearning2024; wangInvestigatingDesigningTrust2024; mohsinCanWeTrust2024), or how LLMs can propagate harmful social biases within their code (liuUncoveringQuantifyingSocial2023; huangBiasTestingMitigation2023) - not whether the coding choices themselves include bias. A study has been conducted on the similarities between human and LLM preferences in task responses (liDissectingHumanLLM2024), but it does not go in depth on the preferences of code-generation.
@@
LLM-based Code Generation: Code-generation in LLMs has been extensively studied (jiangSurveyLargeLanguage2024; zhangUnifyingPerspectivesNLP2023), yet the best way to evaluate model performance on code-generation tasks is an open problem (paulBenchmarksMetricsEvaluations2024a). Existing approaches typically use benchmarks that contain a dataset of programming problems and tests to check the functionality of the generated code (paulBenchmarksMetricsEvaluations2024a), although early benchmarks have shown to be limited in their scope (yadavPythonSagaRedefiningBenchmark2024a). Leveraging LLMs to perform there own code-evaluation has also been suggested (tongCodeJudgeEvaluatingCode2024), as well as the need to test LLMs in more realistic, repository-level, code-generation scenarios (jimenezSWEbenchCanLanguage2024a).
@@
Library-oriented Code: Using external libraries to enhance one’s code is a key part of modern software development (somervilleSoftwareEngineeringGlobal2016). There are multiple studies that discuss the generation of library-oriented code (zanCERTContinualPretraining2022; liuCodeGen4LibsTwoStageApproach2023), including when private-libraries need to be used (zanPrivateLibraryOrientedCodeGeneration2023), and the API-recommendation problem (maCompositionalAPIRecommendation2024; wangSystematicEvaluationLarge2024) within LLMs. But none that discuss the impact that an LLMs preferences towards libraries may have on the coding landscape.
@@
LLMs for Software Engineering: This paper argues that LLMs will impact how a developer makes decisions during software development. There are papers that discuss the problems of library selection (tanzilHowPeopleDecide2024; larios-vargasSelectingThirdpartyLibraries2020), and making implementation design choices (liangQualitativeStudyImplementation2023) by developers, without the use of LLMs. Using ChatGPT as a software librarian to aid these decisions has been suggested (latendresseChatGPTGoodSoftware2024), as well as how users’ distrust toward LLMs may impact the choices they make when using them as coding assistants (choudhuriWhatGuidesOur2024). An empirical study has also been done to analyse how an LLM can best benefit a software engineering project (rasnayakaEmpiricalStudyUsage2024).
@@ 3. Experimental Design
This section sets out the research questions (RQs), describing the methodology used and experiments designed to answer them.
@@ 1. RQ1:
    Language Preference: What programming language preferences do LLMs exhibit in code generation? To answer this RQ, we analyse programming language preference in LLMs using widely-used benchmarks (Section 3.3.1) and more realistic project initialisation tasks (Section 3.3.2).
@@ 2. RQ2:
    Library Preference: What library preferences do LLMs exhibit in code generation? To answer this RQ, we analyse coding library preference in LLMs using library-agnostic Python problems from BigCodeBench (Section 3.4.1) and more realistic project initialisation tasks (Section 3.4.2).
@@ 3. RQ3:
    Internal Consistency: What is the internal consistency of LLMs between the programming languages and libraries they recommend for a task and those they actually use? To answer this RQ, we prompt LLMs to rank their recommended programming languages (Section 3.3.2) and Python libraries (Section 3.4.2). We then evaluate the alignment between their recommendations and actual choices using Kendall’s $\tau$ coefficient.
@@ 4. RQ4:
    LLM Similarities: Do different LLMs have similar preferences when selecting programming languages and libraries? To answer this RQ, we examine the most frequently used languages and libraries across different LLMs in all experiments (Sections 3.3.13.3.23.4.13.4.2) and assess their similarity using Kendall’s $\tau$ coefficient to measure rank correlation across LLMs.
@@ 3.2. LLM Selection
A wide-range of LLMs are used for this study, to gain a broad understanding of the preferences of LLMs when generating code across models of different types. LLMs are chosen to vary in number of parameters, availability (open or closed source), and intended use-case (either general or code-specific). Eight LLMs were chosen, with full details given in Table 1.
We do not use Code Llama in this study. Although it is a popular LLM for code-generation, the markdown text it provides as output does not indicate the programming language being used, as described in Section 3.5.1. Therefore the automated extraction of languages from responses could not be performed reliably.
Model Usage: Each LLM is prompted with default settings and zero-shot prompting to reflect typical developer use. All prompts are run in fresh sessions to avoid bias from caching or leakage.
@@
Model Version Release Knowledge cut-off Size Open- source? Code model? GPT-4o (openaiGPT4TechnicalReport2024) gpt-4o-mini- 2024-07-18 July ’24 Oct. 2023 - GPT-3.5 (andrewpengGPT35TurboFinetuning) gpt-3.5-turbo- 0125 Nov. ’22 Sep. ’21 - Sonnet3.5 (anthropicClaude3Model2024) claude-3-5- sonnet-20241022 Oct. 2024 July 2024 - Haiku3.5 (anthropicClaude3Model2024) claude-3-5- haiku-20241022 Oct. ’24 July ’24 - Llama3.2 (grattafioriLlama3Herd2024) llama-3.2-3b- instruct-turbo Sep. ’24 Dec. ’23 3B Qwen2.5 (huiQwen25CoderTechnicalReport2024) qwen2.5-coder- 32b-instruct Nov. ’24 Mar. ’24 32B DeepSeekLLM (deepseek-aiDeepSeekLLMScaling2024) deepseek-llm- 67b-chat Nov. ’23 May ’23 67B Mistral7B (jiangMistral7B2023) mistral-7b- instruct-v0.3 May ’24 May ’24 7B
3.3. Experimental Setup for Studying Language Preferences

### v2

# A Study of LLMs’ Preferences for Libraries and Programming Languages

@@ Abstract.
Large Language Models (LLMs) are increasingly used to generate code, influencing users’ choices of libraries and programming languages in critical real-world projects. However, little is known about their systematic biases or preferences toward certain libraries and programming languages, which can significantly impact software development practices. To fill this gap, we perform the first empirical study of LLMs’ preferences for libraries and programming languages when generating code, covering eight diverse LLMs. Our results reveal that LLMs exhibit a strong tendency to overuse widely adopted libraries such as NumPy; in up to 48% of cases, this usage is unnecessary and deviates from the ground-truth solutions. LLMs also exhibit a significant preference toward Python as their default language. For high-performance project initialisation tasks where Python is not the optimal language, it remains the dominant choice in 58% of cases, and Rust is not used a single time. These results indicate that LLMs may prioritise familiarity and popularity over suitability and task-specific optimality. This will introduce security vulnerabilities and technical debt, and limit exposure to newly developed, better-suited tools and languages. Understanding and addressing these biases is essential for the responsible integration of LLMs into software development workflows.
@@ 1. Introduction
The last few years have seen Large Language Models (LLMs) make huge advances (naveedComprehensiveOverviewLarge2024a), particularly excelling in code generation tasks (jiangSurveyLargeLanguage2024). Widely accessible tools–such as ChatGPT (openaiGPT4TechnicalReport2024) and GitHub Copilot (chenEvaluatingLargeLanguage2021)–enable developers of all experience levels to use generative AI in their workflow, leading to a significant impact on newly produced code (staffSurveyAIWave2024; lanzStabilityAICEO2023). Existing research has focused mainly on evaluating and improving LLM correctness, security, and efficiency in generating code (chenSurveyEvaluatingLarge2024). Although LLMs demonstrate impressive performance, little is known about their preferences when selecting libraries and programming languages for various coding tasks.
This gap is particularly significant as LLMs are increasingly used as coding assistants by both professional developers and by end users who often lack the expertise to assess whether a recommended library or programming language is appropriate (akhorozConversationalAICoding2025; khuranaWhyWhenLLMBased2024). If LLMs exhibit systematic biases in libraries and programming languages, they can mislead users into making flawed decisions, compromising the reliability, security, and maintainability of software.
@@
To fill this gap, this paper provides the first empirical investigation of the library and programming language preferences of LLMs when generating code. To ensure a broad understanding of LLM preferences, eight widely-used LLMs are chosen for the study, covering both open- and closed-source and a range of sizes: GPT-4o (openaiGPT4TechnicalReport2024), GPT-3.5 (andrewpengGPT35TurboFinetuning), Claude3.5 (Haiku and Sonnet) (anthropicClaude3Model2024), Llama3.2 (grattafioriLlama3Herd2024), Qwen2.5 (huiQwen25CoderTechnicalReport2024), DeepSeekLLM (deepseek-aiDeepSeekLLMScaling2024) and Mistral7b (jiangMistral7B2023).
We begin with libraries–vital components of modern software development (somervilleSoftwareEngineeringGlobal2016)–because developers often omit them from prompts (haoEmpiricalStudyDevelopers2024) and may not know which dependencies a task requires (larios-vargasSelectingThirdpartyLibraries2020), giving LLMs the opportunity to influence library adoption. We then consider the programming language preferences of LLMs. Although professional developers may specify a programming language in the prompt, omitting this information allows us to probe the inherent biases LLMs exhibit when left to make coding choices on their own. This analysis is particularly important as more end users, especially those with limited programming experience, increasingly rely on LLMs’ default choices, often lacking the expertise to evaluate which languages are most appropriate for a given task (nguyenHowBeginningProgrammers2024), which is a trend especially evident in “vibe-coding” scenarios (smithAIVibeCoding2025).
@@
For both libraries and programming languages, we examine LLM preferences in two key scenarios: writing code for practical tasks from widely studied benchmarks; and generating the initial code for new projects, a crucial phase where LLMs have been found out to be particularly helpful (rasnayakaEmpiricalStudyUsage2024) and foundational technology choices are made. Additionally, we explore whether the coding technologies that LLMs provide as recommendations match those they actually use when generating code, and whether all studied LLMs share similar preferences.
@@
Our experiments lead to the following primary observations: (1) All LLMs we study heavily favour well established libraries over high-quality alternatives. In particular, LLMs tend to overuse widely adopted data science libraries. For example, NumPy is used unnecessarily in up to 48% of cases, diverging from the ground-truth solutions. (2) For programming languages, all LLMs we study exhibit a significant preference toward using Python. To our surprise, even for tasks where high performance and memory safety are critical–and therefore Python is suboptimal–Python remains the dominant choice as the most used language in 58% of cases, while Rust is not used a single time. (3) There is extremely low consistency between the libraries and programming languages recommended by LLMs and those they actually use.
LLMs may develop these strong biases due to various factors, such as the prevalence of certain libraries in open source repositories, the distribution of programming languages included in their training data (wangExploringMultiLingualBias2024), or post-training alignment and fine-tuning processes (zieglerFineTuningLanguageModels2020). For example, if an LLM is trained on codebases where a particular library is favoured, it may exhibit a strong inclination toward generating code that uses the given library, even in cases where it might not be necessary.
Understanding and mitigating programming language and library bias in LLMs is not merely an academic concern. The systemic biases we observe have immediate and far-reaching practical implications: (1) Strong biases in default choices introduce technical debt and security risks because of suboptimal choices, this is especially problematic for safety-critical systems. (2) LLMs now frequently augment training data with self-generated data (zhuHowSynthesizeText2025), where language and library favouritism will create a dangerous feedback loop, further reducing diversity in synthetic data. (3) Bias toward well established languages and libraries actively blocks newer languages and open-source tools. This not only creates a discoverability barrier for newer or community-driven tools but also inhibits innovation, particularly in open-source environments.
We call for contributions from the research and developer communities to better understand, evaluate, and mitigate these biases, ensuring that LLMs support a more diverse, responsible, and forward-looking software development landscape. The primary contributions of this paper are as follows.
@@
  • Presenting the first empirical study of LLM preferences for libraries and programming languages when generating code, revealing a clear systemic bias and lack diversity, with significant preferences toward established libraries and Python.
@@
  • Increasing awareness of the need to measure and understand the diversity and preferences within LLM-generated code, providing a discussion of how this could impact the coding landscape and open-source software ecosystem.
@@
  • Releasing our code, datasets and complete results publicly via our GitHub repository [^1] Project repository: https://github.com/itsluketwist/llm-code-bias, to encourage further investigation in this area.
@@
The rest of the paper is organised as follows. Section 2 summarises work in related areas. Section 3 details the research questions, and the experiments conducted to answer them. Section 4 presents the results and answers to the research questions. Section 5 further discusses the findings and their broader impacts. Section 6 presents the threats to validity. Section 7 concludes the paper.
@@ 2. Related Work
To our knowledge, this is the first study to examine LLM biases for libraries and programming languages when writing code. In the following, we cover related questions from prior work.
@@
Bias and Unfairness in LLMs: Whether an LLM has preferences for specific programming languages and libraries is related to bias and whether an LLM provides an unfair representation of the world. Bias within an LLM’s natural language (NL) responses has been extensively studied (gallegosBiasFairnessLarge2024), but bias within any generated code is widely under explored. Most works focus on the trustworthiness of the code (patelEvaluatingInContextLearning2024; wangInvestigatingDesigningTrust2024; mohsinCanWeTrust2024), or how LLMs can propagate harmful social biases within their code (liuUncoveringQuantifyingSocial2023; huangBiasTestingMitigation2023) - not whether the coding choices themselves include bias. A study has been conducted on the similarities between human and LLM preferences in task responses (liDissectingHumanLLM2024), but it does not go into depth on code generation preferences.
@@
LLM-based Code Generation: Code-generation in LLMs has been extensively studied (jiangSurveyLargeLanguage2024; zhangUnifyingPerspectivesNLP2023), yet the best way to evaluate the performance of LLMs on code generation tasks is an open problem (paulBenchmarksMetricsEvaluations2024a). Existing approaches typically use benchmarks that contain a dataset of programming problems and tests to check the functionality of generated code (paulBenchmarksMetricsEvaluations2024a), although early benchmarks have been shown to be limited in their scope (yadavPythonSagaRedefiningBenchmark2024a). Using LLMs to perform their own code evaluation has also been suggested (tongCodeJudgeEvaluatingCode2024), as well as the need to test LLMs in more realistic repository-level code-generation scenarios (jimenezSWEbenchCanLanguage2024a).
@@
Library-oriented Code: Using libraries to enhance code is a key part of modern software development (somervilleSoftwareEngineeringGlobal2016). Multiple studies discuss the generation of library-oriented code (zanCERTContinualPretraining2022; liuCodeGen4LibsTwoStageApproach2023), including when private libraries must be used (zanPrivateLibraryOrientedCodeGeneration2023), and the API recommendation problem (maCompositionalAPIRecommendation2024; wangSystematicEvaluationLarge2024) within LLMs. None of the above discuss the potential impact of bias in LLM library preferences.
@@
Software Engineering Decisions: This paper argues that LLMs will affect how a developer makes decisions during software development. The literature discusses how developers select libraries (tanzilHowPeopleDecide2024; larios-vargasSelectingThirdpartyLibraries2020), and make implementation choices (liangQualitativeStudyImplementation2023), without the use of LLMs. Using ChatGPT as a software librarian has been suggested to aid these decisions (latendresseChatGPTGoodSoftware2024), as well as how users’ distrust towards LLMs may impact the choices they make when using them as coding assistants (choudhuriWhatGuidesOur2024). An study has also been done to analyse how an LLM can best benefit a software engineering project (rasnayakaEmpiricalStudyUsage2024).
@@ 3. Experimental Design
This section presents the research questions (RQs) and describes the experiments designed to answer them. For transparency and reproducibility, we have released the full code for all experiments described here on our GitHub repositoryfootnotemark: .
@@ 1. RQ1:
    Library Preference: What library preferences do LLMs exhibit in code generation? To answer this RQ, we analyse the library preferences in LLMs using library-agnostic benchmark tasks (Section 3.3.1) and more realistic project initialisation tasks (Section 3.3.2).
@@ 2. RQ2:
    Language Preference: What programming language preferences do LLMs exhibit in code generation? To answer this RQ, we analyse the programming language preferences in LLMs using widely adopted benchmark tasks (Section 3.4.1) and more realistic project initialisation tasks (Section 3.4.2).
@@ 3. RQ3:
    Recommendation Consistency: Is there any consistency between the libraries and programming languages LLMs recommend for a task and those they actually use? To answer this RQ, we ask LLMs to rank their recommended libraries (Section 3.3.2) and programming languages (Section 3.4.2) for project initialisation tasks; then evaluate the alignment between their recommendations and actual choices using Kendall’s $\tau$ coefficient.
@@ 4. RQ4:
    LLM Similarities: Do different LLMs have similar preferences when selecting libraries and programming languages? To answer this RQ, we examine the libraries and programming languages used in all experiments (Sections 3.3.1, 3.3.2, 3.4.1, 3.4.2) and assess the similarity between LLMs using Kendall’s $\tau$ coefficient.
@@ 3.2. LLM Selection
We use a diverse set of LLMs in this study, to gain a broad understanding of LLM preferences. LLMs are chosen to vary in number of parameters, availability (open- or closed-source), and intended use case (general or code-specific). This setup enables us to explore whether the underlying architecture influences the results. Eight LLMs were chosen, with full details given in Table 1.
LLM Usage: To reflect the typical usage of LLMs by developers, which often overlooks the role of parameters (donatoStudyingHowConfigurations2025), each LLM is prompted using the default values for the temperature and top_p parameters. Closed-source LLMs are prompted using the default values of their corresponding APIs; open-source models are prompted using the default values given on each LLM’s Hugging Face [^2] Hugging Face: https://huggingface.co/models repository. The parameter values are given in Table 1.
Furthermore, we conduct each LLM interaction in a fresh API session, to avoid bias from prompt caching or leakage (guAuditingPromptCaching2025); and we do not use a system prompt to ensure that each LLM has its base functionality considered without external influence (muCloserLookSystem2025).
@@
Model Version Release Knowledge cut-off Open- source? Code model? Size Parameters temp top_p GPT-4o (openaiGPT4TechnicalReport2024) gpt-4o-mini- 2024-07-18 July ’24 Oct. ’23 - 1.0 1.0 GPT-3.5 (andrewpengGPT35TurboFinetuning) gpt-3.5-turbo- 0125 Nov. ’22 Sep. ’21 - 1.0 1.0 Sonnet3.5 (anthropicClaude3Model2024) claude-3-5- sonnet-20241022 Oct. ’24 July ’24 - 1.0 1.0 Haiku3.5 (anthropicClaude3Model2024) claude-3-5- haiku-20241022 Oct. ’24 July ’24 - 1.0 1.0 Llama3.2 (grattafioriLlama3Herd2024) llama-3.2-3b- instruct-turbo Sep. ’24 Dec. ’23 3B 0.6 0.9 Mistral7B (jiangMistral7B2023) mistral-7b- instruct-v0.3 May ’24 May ’24 7B 0.7 1.0 Qwen2.5 (huiQwen25CoderTechnicalReport2024) qwen2.5-coder- 32b-instruct Nov. ’24 Mar. ’24 32B 0.7 0.8 DeepSeekLLM (deepseek-aiDeepSeekLLMScaling2024) deepseek-llm- 67b-chat Nov. ’23 May ’23 67B 0.7 0.95
3.3. Experimental Setup for Library Preferences
To investigate library preferences, we focus on a single programming language to enable a more in-depth analysis. Python is chosen for library-related experiments due to its vast collection of open-source libraries [^3] Python Package Index: https://pypi.org/, its easy-to-parse import syntax (as discussed in 3.5.1), and its popularity in the open-source community (staffOctoverseAILeads2024a).
Previous work shows that users often start tasks with NL descriptions without knowing which–or that any–libraries are required (kuhnDesigningBetterTools2012); and when prompting LLMs, will initially omit details, iteratively refining their requirements over time (haoEmpiricalStudyDevelopers2024). Therefore, we investigate library preferences using tasks that omit the specific libraries to use, to mirror real‑world prompting.